# PolyOCP.jl– A Julia Package for Stochastic OCPs and MPC

### v2

@@ II-C Stochastic optimal control
The data structure for defining (10) is a struct called StochProb whose fields are described in Table III. The Boolean flag gauss determines how $\gamma(\varepsilon)$ is selected in the chance constraint (10c) as discussed following (10). Note that chance constraints in PolyOCP.jl are defined on one side as $\mathbb{P}[Z\leq\texttt{bound}]\geq 1-\texttt{risk}$, where setting bound=Inf effectively deactivates the constraint.
Given the required and optional fields, the function build automatically generates the optimization problem corresponding to  (10) using a user-specified solver, e.g.Ipopt, through JuMP.jl, which is a domain-specific model language for expressing and solving mathematical optimization problems [24, 37]. Finally, the function solveOCP solves  (10) and returns the optimal solutions in PCE coefficients $\textsf{x}^{j,\star}$ and $\textsf{u}^{j,\star}$ for $j\in\mathbb{I}_{[0,L-1]}$.
For using PolyOCP.jl in Model Predictive Control (MPC), the function con_initial_param sets up a constraint for the initial condition with the parametric variable x0_Param to avoid rebuilding the entire model repeatedly. The value of x0_Param is then updated at each time step with the current measured state using update_initial_param in the MPC loop.
TABLE III: Structure StochProb defining OCP (10)
@@
Except for defining (10) through the provided struct StochProb, PolyOCP.jl also offers additional functions to enhance modeling flexibility.
When objectives beyond quadratic costs—such as risk-aware objectives—are considered, PolyOCP.jl allows users to specify custom objective functions by leaving the weighting matrix fields in StochProb empty. If no objective is defined, JuMP.jl typically returns a feasible solution.
When only a subset of state or input components is subject to chance constraints, the fields lbx, ubx, lbu, and ubu may be omitted. Instead, chance constraints can be conveniently imposed using the function con_chance that provides an interface for specifying individual constraints. Furthermore, after building the optimization problem via build, additional constraints can also be appended following the standard syntax of JuMP.jl.
To compute the Probability Density Function (PDF) of a random variable from its PCE $Z=\sum_{j=0}^{L-1}\textsf{z}^{j}\psi^{j}(\xi)\in\mathcal{L}^{2}(\mathbb{R})$ numerically, one can employ the Fourier transformation $\mathcal{F}$ and its inverse $\mathcal{F}^{-1}$ as
@@
Especially, the Fourier transforms of the canonical distributions listed in Table II are known as characteristic functions [38].
@@ IV-A Chemical reactor – Stochastic OCP
$A=\begin{bmatrix}0.95123&0\\ 0.08833&0.81873\end{bmatrix},\,B=\begin{bmatrix}-0.0048771\\ -0.0020429\end{bmatrix},\,E=\begin{bmatrix}1\\ 1\end{bmatrix}$.
The initial condition $X_{\text{ini}}$ follows a 2-variate Gaussian distribution as $X_{\text{ini}}\sim\mathcal{N}([0.5,\,0.1]^{\top},\mathrm{Diag}([0.05^{2},\,0.01^{2}]))$, and disturbances $W$ are uniformly distributed on [0.0173,0.0173][-0.0173,0.0173]. As suggested in Table II, a 2-variate Hermite polynomials $\{\psi(\xi_{\text{ini}})\}_{i=0}^{2}=\{1,\xi_{\text{ini},1},\xi_{\text{ini},2}\}$ is used for $X_{\text{ini}}$, where $\xi_{\text{ini},1}$ and $\xi_{\text{ini},2}$ are independent standard Gaussian variables. For $W_{k}$, $k\in\mathbb{I}_{[0,N-1]}$, Jacobi polynomials $\{\psi_{w}\}_{n=0}^{1}=(\xi_{k})=\{1,2\sqrt{3}(\xi_{k}-0.5)\}$ are employed, where $\xi_{k}$ are i.i.d. variables following $\mathcal{U}(0,1)$. Their PCEs and the joint basis from (8) are then given by
$X_{\text{ini}}$ $=\begin{bmatrix}0.5\\ 0.1\end{bmatrix}\cdot 1+\begin{bmatrix}0.05\\ 0\end{bmatrix}\cdot\psi^{1}(\xi_{\text{ini}})+\begin{bmatrix}0\\ 0.01\end{bmatrix}\cdot\psi^{2}(\xi_{\text{ini}}),$
$W_{k}$ $=0\cdot 1+\frac{2\cdot 0.0173}{2\sqrt{3}}\psi^{1}(\xi_{k}),\quad k\in\mathbb{I}_{[0,N-1]},$
@@ IV-B Four tank system – Stochastic MPC
$A=\begin{bmatrix}[l]0.921&0&0.041&0\\ 0&0.918&0&0.033\\ 0&0&0.924&0\\ 0&0&0&0.937\end{bmatrix},\,B=\begin{bmatrix}[l]0.017&0.001\\ 0.001&0.023\\ 0&0.061\\ 0.072&0\end{bmatrix}$,
@@
where $\psi^{1}(\xi)=\xi$ and $\psi^{2}=(\xi^{2}-1)/\sqrt{2}$ are listed in Table II. Each component of the initial condition $X_{\text{ini}}$ is uniformly distributed on [1,1][-1,1]. Moreover, the weighting matrices are chosen to be $Q=Q_{N}=3\cdot I_{4}$ and $R=10^{4}\cdot I_{2}$. The chance constraints are imposed on the first two components of the state as $\mathbb{P}[-2\leq X_{i}\leq 2]\geq 0.8$, $i=1,2$. Similar to Example IV-A,the corresponding StochProb with prediction horizon $N=10$ is constructed and the model is built. Note that when running PolyOCP.jl in MPC, the parametric initial condition x0Param is set to the current measured state. Further details about this example are provided in examples/Tank.jl.
@@
Further examples, e.g., stochastic LTI systems with non-i.i.d. addictive disturbances, are available under /examples.
@@ V Conclusion
This paper has introduced PolyOCP.jl, a julia toolbox, released under the MIT license, for solving stochastic optimal control problems in the PCE framework. Existing PCE tools are not tailored to handle problems with additive stochastic disturbances. PolyOCP.jl presents a first step towards closing this gap. The toolbox is intended to facilitate rapid prototyping of stochastic OCPs and corresponding MPC formulations.

### v3

@@ II-C Stochastic optimal control
The data structure for defining (10) is a struct called StochOCP whose fields are described in Table III. The Boolean flag gauss determines how $\gamma(\varepsilon)$ is selected in the chance constraint (10c) as discussed following (10). Note that chance constraints in PolyOCP.jl are defined on one side as $\mathbb{P}[Z\leq\texttt{bound}]\geq 1-\texttt{risk}$, where setting bound=Inf effectively deactivates the constraint.
Given the required and optional fields, the function buildOCP automatically generates the optimization problem corresponding to  (10) using a user-specified solver, e.g. Ipopt, through JuMP.jl, which is a domain-specific model language for expressing and solving mathematical optimization problems [24, 37]. Finally, the function solveOCP solves  (10) and returns the optimal solutions in PCE coefficients $\textsf{x}^{j,\star}$ and $\textsf{u}^{j,\star}$ for $j\in\mathbb{I}_{[0,L-1]}$.
For using PolyOCP.jl in Model Predictive Control (MPC), the function con_initial_param sets up a constraint for the initial condition with the parametric variable x0param to avoid rebuilding the entire model repeatedly. The value of x0param is then updated at each time step with the current measured state using update_initial_param in the MPC loop.
TABLE III: Structure StochOCP defining OCP (10)
@@
Except for defining (10) through the provided struct StochOCP, PolyOCP.jl also offers additional functions to enhance modeling flexibility.
When objectives beyond quadratic costs—such as risk-aware objectives—are considered, PolyOCP.jl allows users to specify custom objective functions by leaving the weighting matrix fields in StochOCP empty. If no objective is defined, JuMP.jl typically returns a feasible solution.
When only a subset of state or input components is subject to chance constraints, the fields lbx, ubx, lbu, and ubu may be omitted. Instead, chance constraints can be conveniently imposed using the function con_chance that provides an interface for specifying individual constraints. Furthermore, after building the optimization problem via buildOCP, additional constraints can also be appended following the standard syntax of JuMP.jl.
To compute the Probability Density Function (PDF) of a random variable from its PCE $Z=\sum_{j=0}^{L-1}\textsf{z}^{j}\psi^{j}(\xi)\in\mathcal{L}^{2}(\mathbb{R})$ numerically, one can use the function pdfPCE that employs the Fourier transformation $\mathcal{F}$ and its inverse $\mathcal{F}^{-1}$ as
@@
The Fourier transforms of the canonical distributions listed in Table II are also known as characteristic functions [38].
@@ IV-A Chemical reactor – Stochastic OCP
$A=\left[\begin{matrix}[l]0.95123&0\\ 0.08833&0.81873\end{matrix}\right],\,B=\left[\begin{smallmatrix}-0.0048771\\ -0.0020429\end{smallmatrix}\right],\,E=\left[\begin{smallmatrix}1\\ 1\end{smallmatrix}\right].$
The initial condition $X_{\text{ini}}$ follows a 2-variate Gaussian distribution as $X_{\text{ini}}\sim\mathcal{N}([0.5,\,0.1]^{\top},\mathrm{Diag}([0.05^{2},\,0.01^{2}]))$, and disturbances $W$ are uniformly distributed on [0.2076, 0.2076][-0.2076,\,0.2076]. As suggested in Table II, a 2-variate Hermite polynomial basis $\{\psi(\xi_{\text{ini}})\}_{i=0}^{2}=\{1,\xi_{\text{ini},1},\xi_{\text{ini},2}\}$ is used for $X_{\text{ini}}$, where $\xi_{\text{ini},1}$ and $\xi_{\text{ini},2}$ are independent standard Gaussian variables. For $W_{k}$, $k\in\mathbb{I}_{[0,N-1]}$, Jacobi polynomials $\{\psi_{w}^{n}(\xi_{k})\}_{n=0}^{1}=\{1,\,2\sqrt{3}(\xi_{k}-0.5)\}$ are employed, where $\xi_{k}$ are i.i.d. variables following $\mathcal{U}(0,1)$. Their PCEs and the joint basis from (8) are then given by
$X_{\text{ini}}$ $=\begin{bmatrix}0.5\\ 0.1\end{bmatrix}\cdot 1+\begin{bmatrix}[l]0.05\\ 0\end{bmatrix}\cdot\psi^{1}(\xi_{\text{ini}})+\begin{bmatrix}[l]0\\ 0.01\end{bmatrix}\cdot\psi^{2}(\xi_{\text{ini}}),$
$W_{k}$ $=0\cdot 1+\frac{2\cdot 0.2076}{2\sqrt{3}}\psi^{1}(\xi_{k}),\quad k\in\mathbb{I}_{[0,N-1]},$
@@ IV-B Four tank system – Stochastic MPC
$A=\left[\begin{matrix}[l]0.921&0&0.041&0\\ 0&0.918&0&0.033\\ 0&0&0.924&0\\ 0&0&0&0.937\end{matrix}\right],\,B=\left[\begin{matrix}[l]0.017&0.001\\ 0.001&0.023\\ 0&0.061\\ 0.072&0\end{matrix}\right],$
@@
where $\psi^{1}(\xi)=\xi$ and $\psi^{2}=(\xi^{2}-1)/\sqrt{2}$ are listed in Table II. Each component of the initial state $X(0)$ is sampled uniformly from [1,1][-1,1] to generate different initial conditions. Moreover, the weighting matrices are chosen to be $Q=Q_{N}=3\cdot I_{4}$ and $R=10^{4}\cdot I_{2}$. The chance constraints are imposed on the first two components of the state as $\mathbb{P}[-2\leq X_{i}\leq 2]\geq 0.8$, $i=1,2$. Similar to Example IV-A, the corresponding StochOCP with prediction horizon $N=10$ is constructed and the model is built. Note that when running PolyOCP.jl in MPC, the parametric initial condition x0param is set to the current measured state. Further details about this example are provided in examples/Tank.jl.
@@
Further examples, e.g., stochastic LTI systems with non-i.i.d. additive disturbances, are available under /examples.
@@ V Conclusion
This paper has introduced PolyOCP.jl, a julia toolbox, released under the MIT license, for solving stochastic optimal control problems in the PCE framework. Existing PCE tools are not tailored to handle problems with additive stochastic disturbances. PolyOCP.jl presents a first step towards closing this gap. The toolbox is intended to facilitate rapid prototyping of stochastic OCPs and corresponding stochastic MPC schemes.